OpenStudy (anonymous):

Solve the system of linear equations below. x + y = 4 2x + 3y = 0

OpenStudy (anonymous):

x+y=4 multiply by-2 & add. 2x+3y=0 -2X-2Y=-8 Y=-8 ans. 2X+3*-8=0

OpenStudy (anonymous):

x = 12